OpenStudy (anonymous):

Solve the following quadratic equation by using the quadratic formula 6^2 - 16x - 6 = 0

OpenStudy (anonymous):

typo somewheres i bet

OpenStudy (anonymous):

\[x^2-16x-6=0\] or maybe \[6x^2-16x-6=0\]is my guess

pooja195 (pooja195):

\[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

pooja195 (pooja195):

find your abc values \[\huge~\rm~ax^2+bx+c=0 \]

OpenStudy (anonymous):

\[6x^2 - 16x - 6 = 0\] Is the equation.

OpenStudy (bee_see):

A is 6, b is -16, and c is -6. Plug in the numbers